You have an equilateral triangle, and it happens you place three charges 45, -35, and 77 on each vertex, the equilateral triangle is separated by the distance 61. What is the total net force experience by -35 due to other two charges if the charge is place at the middle vertex of the triangle?



1
Expert's answer
2021-01-14T16:35:21-0500

Let the first and third charges located at the base of the equilateral triangle be "45\ C" and "77\ C", respectively. Let the second charge located at the middle vertex of the equilateral triangle be "-35\ C". Let, also, the length of the side of the equilateral triangle be "r=61\cdot10^{-15}\ m" and the side "q_1q_2" be the "x"-axis.

The force "F_{12}" directed away from the positive charge "q_1"(toward the negative charge "q_2"). The force "F_{32}" directed away from the positive charge "q_3"(toward the negative charge "q_2"). It is obviously, that the net electric force on charge "q_2" due to charges "q_1" and "q_3" is the vector sum of forces "F_{12}" and "F_{32}".

Let’s find the magnitudes of these forces:


"F_{12}=k\dfrac{|q_1q_2|}{r^2},""F_{12}=9\cdot10^9\ \dfrac{Nm^2}{C^2}\dfrac{|45\ C\cdot(-35\ C)|}{(61\cdot10^{-15}\ m)^2}=3.81\cdot10^{39}\ N,""F_{32}=k\dfrac{|q_3q_2|}{r^2},""F_{32}=9\cdot10^9\ \dfrac{Nm^2}{C^2}\dfrac{|77\ C\cdot(-35\ C)|}{(61\cdot10^{-15}\ m)^2}=6.52\cdot10^{39}\ N."


Then, we can find the projections of forces "F_{12}" and "F_{32}" on axis "x" and "y":


"F_x=F_{12x}+F_{32x},""F_x=3.81\cdot10^{39}\ N+6.52\cdot10^{39}\ N\cdot cos60^{\circ}=7.07\cdot10^{39}\ N,""F_y=F_{12y}+F_{32y},""F_y=0+6.52\cdot10^{39}\ N\cdot sin60^{\circ}=5.65\cdot10^{39}\ N."

Finally, the net electric force will be:


"F_{net}=\sqrt{F_x^2+F_y^2},""F_{net}=\sqrt{(7.07\cdot10^{39}\ N)^2+(5.65\cdot10^{39}\ N)^2}=9.05\cdot10^{39}\ N."

Answer:

"F_{net}=9.05\cdot10^{39}\ N."